A bus arrives every 10 minutes at a bus stop. It

is assumed that the waiting time for a particular individual is a random variable with a continuous uniform

distribution.(a) What is the probability that the individual waits

more than 7 minutes?

(b) What is the probability that the individual waits

between 2 and 7 minutes?


1
Expert's answer
2022-01-11T10:36:07-0500

(a) "P(X>7)=\frac{10-7}{10}=0.3."


(b) "P(2<X<7)=\frac{7-2}{10}=0.5."